you begin with $90 in your savings account and your friend begins with $35 in her savings account. You deposit $10 in savings each week, and your friend deposits $15 in savings each week. write and graph a system of linear equations that represent the amounts in each of your saving accounts. Your friend says that in 10 weeks you will both have the same amount of money in your savings accounts. is your friend correct? use the graph from part a to explain your answer.
Answer by Danielshinmath(31)

You can represent your total money as y= 10x+90 and your friends as y=15x+35, where your 10 and your friends 15 mean how much deposited each week. the x is the amount of weeks, and the last number is how much is already stored- the 90 and the 35. The y is the total money. Your friend said that you will have the same amount as her in 10 weeks. Let's see if that is true. input 10 for both X's and since they both equal y, you can do:
10x+90=15x+35.---> 10(10)+90=15(10)+35---->190=185. This is not true, so your friend is wrong.